OpenStudy (johnt):

Determine whether the series converges or diverges. 3^(n+2) / 5^n

OpenStudy (johnt):

\[a_n = \frac{ 3^{n+2} }{ 5^n } \]

OpenStudy (johnt):

I'm not sure where to start with this. WolframAlpha says the limit is 0, suggesting it converges, but I don't know how to prove that. It looks like you'd use L'Hospital's rule, but that doesn't do anything. You just get \[3^{n+2} \ln(3)\] again, which doesn't help to integrate it. If someone could help me understand where to begin, I'd appreciate it.

OpenStudy (e.mccormick):

Hmmm.... yah, that one is a little past me, but since nobody posted for an hour, let me link something that might cover it: http://tutorial.math.lamar.edu/Classes/CalcII/ConvergenceOfSeries.aspx

OpenStudy (anonymous):

Just because \(\displaystyle\lim_{n\to\infty}a_n=0\) doesn't mean \(\sum a_n\) converges. The converse is true, though: If \(\sum a_n\) converges, then \(\displaystyle\lim_{n\to\infty}a_n=0\). For this series, you just have to do some rewriting: \[\begin{align*}\sum_{n=0}^\infty \frac{3^{n+2}}{5^n}&=\sum_{n=0}^\infty \frac{3^{n}\cdot3^2}{5^n}\\ &=9\sum_{n=0}^\infty \frac{3^n}{5^n}\\ &=9\sum_{n=0}^\infty \left(\frac{3}{5}\right)^n\end{align*}\] What do you know about the convergence/divergence of a geometric series?